OpenStudy (anonymous):

What is the value of x to the nearest degree? A. 32 B. 58 C. 79 D. 83

OpenStudy (anonymous):



OpenStudy (souvik):

 tanx=AB/BC=8/5 x=tan^-1(8/5)

OpenStudy (anonymous):

i got like 58

OpenStudy (anonymous):

OH. so the answer is b?

OpenStudy (souvik):

right!

OpenStudy (anonymous):

thank you so much!

OpenStudy (souvik):

welcome:)